MEASUREMENT OF TORSIONALLY EXCITED MICROWAVE TRANSITIONS IN METHANOL-D_{4}

\underline{INDRANATH MUKHOPADHYAY}, Laser Programme, Centre for Advanced Technology, Indore 452 013, India; K.V.L.N. SASTRY, Department of Physics, University of New Brunswick, Fredericton, N.B., E3B 5A3, Canada; MANFRED WINNEWISSER, Physikalisch Chemisches Institut der Justus Liebig Universitaet, Heinrich-Buff-Ring 58, D-35392 Giessen, Germany.

In this work, 14 new microwave transitions have been measured in CD_{3}OD species of methanol with an estimated accuracy of 100 kHz. The transitions have been assigned to definite torsional-rotational quantum states in the vibrational ground state. The majority of the assigned transitions involve the first excited torsional state. The present data along with the previously known microwave transitions and recently assigned Fourier transform transitions have been fitted with a recently developed Hamiltonian model and the torsion-rotation molecular parameters determined.